Answer the following problem. Show your solution.


1. A stone which has a mass of 1.27 kg is tied to a rope of 9 m and moved in a circular motion. If it is moving at the rate of 5 m/s.What is the centripetal force acting on it?


2. A 1500kg truck, running at a speed of 10.5 m/s.What is the cemtripetal force acting on it?


3. What is the centripetal force of an 9000-kg train, running at a speed of 27.8 m/s and rounding a curve with a curvature radius of 150 m?


1
Expert's answer
2021-09-03T12:12:21-0400

1

"F=\frac{mv^2}{r}=\frac{1.27\:\rm kg*(5\: m/s)^2}{9\:\rm m}=3.5\:\rm N"


2. When the trajectory of motion is a straight line, the centripetal force is zero.

3.

"F=\frac{mv^2}{r}=\frac{9000\:\rm kg*(27.8\: m/s)^2}{150\:\rm m}=46370\:\rm N"